### v2

# Unified Fourier bases for signals on random graphs with group symmetries

Abstract
We consider a recently proposed approach to graph signal processing (GSP) based on graphons. We show how the graphon-based approach to GSP applies to graphs sampled from a stochastic block model derived from a weighted Cayley graph. When SBM block sizes are equal, a nice Fourier basis can be derived from the representation theory of the underlying group. We explore how the SBM Fourier basis is affected when block sizes are not uniform. When block sizes are nearly uniform, we demonstrate that the group Fourier basis closely approximates the SBM Fourier basis. More specifically, we quantify the approximation error using matrix perturbation theory. When block sizes are highly non-uniform, the group-based Fourier basis can no longer be used. However, we show that partial information regarding the SBM Fourier basis can still be obtained from the underlying group.
keywords:
graphon, stochastic block model, graph signal processing, graph/graphon Fourier transform.
{AMS}
Primary 94A12. Secondary 05C50.
1 Introduction
Graph Signal Processing (GSP) has become a popular method to process data described on irregular domains. Such data can often be interpreted as a signal that assigns numerical values to the nodes of a network or graph. GSP has been successfully used in applications such as network failure assessment [29], detection of false data attacks in smart grids [8], and analysis of brain signals [13, 22, 23, 24]. For an overview of GSP and its applications, we refer to [17, 27, 28, 34, 35, 36].
One of the most fundamental concepts in classical signal processing is the Fourier transform. This notion has been generalized to graph signals by using spectral features of the underlying graph. In this approach, we first assign a shift operator to the underlying graph of a signal. The graph Fourier transform is then defined as the projection of signals onto a fixed eigenbasis of the graph shift operator. Using the graph Fourier transform, important concepts of signal processing have been generalized to the case of graph signals.
This spectral approach to GSP has two major drawbacks: firstly, calculating eigenbases for large matrices is computationally expensive and slow; secondly, the graph Fourier basis depends rigidly on the specific graph at hand. However, the underlying graph of a signal may sustain minor variations due to error or changes in the network over time. In this paper, we use the theory of graph limits to provide an instance-independent Fourier transform for samples of stochastic block models, where the Fourier basis is computed according to the common structure of the graphs in a family rather than each particular instance. This approach, which was first proposed in [25], provides several practical benefits. Most importantly, using this method, the problem of finding an eigenbasis for each particular network is reduced to a single computation for the limiting object. In addition, this approach allows us to compare signals on a collection of different graphs with similar high-scale structures (e.g., brain networks of different individuals). A key advantage of the graphon-based approach is its ability to capture approximate large-scale symmetries within a family of graphs. That is, graphs may be modeled as random samples from a limiting graphon that manifests the symmetries, and a Fourier basis for the graphon can be used as a shared basis for the sampled graphs. For example, if the graphon has a circulant structure, then it exhibits the symmetries of a cyclic group. Its cyclic symmetries resemble periodicity of signals in the classical setting, and its Fourier transform closely resembles the traditional discrete Fourier transform. However, graphs sampled from this graphon generally retain the cyclic symmetries only in approximation, and a Fourier basis derived directly from an individual graph may lack distinctive properties. In contrast, the graphon-based approach yields a well-structured Fourier basis that can be effectively applied to any graph sampled from the graphon.
The limit theory of (dense) graph sequences was initiated by Lovász and Szegedy in [19]. Graphs with similar high-scale structures (i.e., similar subgraph densities) are close in a metric derived from the cut norm [18]. Any sequence of graphs, that is Cauchy in cut norm, converges to a limit, and such limits are called graphons. A graphon is a symmetric measurable function on $[0,1]^{2}$ with values in $[0,1]$. The space of graphons includes all labeled graphs, and the definition of cut norm extends to graphons in a natural way. Every graphon $w:[0,1]^{2}\to[0,1]$ gives rise to a rather general random graph model, called the $w$-random graph, whose samples are graphs of any desired size. The $w$-random model provides an excellent sampling scheme: almost every growing sequence generated by a $w$-random model converges to $w$ itself.
Here, we focus our attention on a special class of graphon models, called the Stochastic Block Model (SBM). In the SBM model, vertices are divided into blocks of prescribed proportion. Edges are added independently, with probability determined by the block membership of each vertex. In a large graph sampled from an SBM, the edge density between two given blocks will be approximately equal to the link probability of those two blocks (e.g., see Fig. 1). The SBM is a natural model for many applications and is widely used in network analysis; see [1] for an overview.
Figure 1: Graphon $w$ representing SBM, shown in the middle , taking value 0.8 on gray and 0.2 on white cells. Sampled graph on 60 vertices (left) and one on 600 vertices (right) are shown. For graphs, the vertices have been placed on $[0,1]$, and edges are represented by black pixels.
The use of graphons in developing a common scheme for signal processing on large networks (i.e. instance-independent GSP schemes) was first proposed in 2020 [31], and the robustness (i.e. convergence) of these methods was proved in [12, 32]. Namely, it was proved that if a sequence of graph signals converges to a graphon signal, the graph Fourier transform can be approximated by the Fourier transform of the limiting graphon signal (see Theorem 2.9 for details). To enable the development of simultaneous GSP methods on samples of stochastic block models, we focus on the adjacency matrix (instead of the Laplacian matrix) as our shift operator. This is akin to the choice of shift operator in the definition of graphon signal processing proposed in [32].
In this paper, we focus on graphons with group theoretic symmetries. Samples of such graphons can be viewed as generalizations or perturbations of Cayley graphs. A Cayley graph has vertices corresponding to elements of a group and edges generated by shifts with elements of an inverse-closed subset of the group. The underlying algebraic structure and highly symmetric nature of Cayley graphs make them a rich category of graphs for various applications. On Cayley graphs, performing GSP compatible with the group Fourier basis has desirable properties, see [3, 11]. For examples of signal processing on Cayley graphs, see [7, 15, 16, 30]. We will demonstrate that for SBM with relaxed Cayley structure, some of these favorable properties are preserved.
1.1 Main contribution
In this paper, we discuss how the graphon Fourier transform can be used to provide suitable instance-independent Fourier transforms for samples of SBM, which we call an SBM-driven Fourier transform for signals on such graphs. In particular, we investigate the case of SBM with group theoretic symmetries. Our contribution is three-fold. Firstly, we show how the problem of computing the eigenspace decomposition of an SBM can be turned into a low-dimensional problem, and how the group Fourier basis can be used to find a Fourier basis for SBM with Cayley structure and uniform block sizes. Secondly, we show that relatively small changes in the block sizes do not affect the obtained eigen-decomposition dramatically. Thirdly, we find a basis for stochastic block models with an underlying Cayley structure, even if the symmetry is broken due to unequal block sizes.
1.2 Organization of the paper
In Section 2, we collect notations and background regarding the stochastic block model, graphons, and graph/graphon Fourier transforms. We end this section by quoting Theorem 2.9, which lays the groundwork for developing instance-independent GSP. In Section 3, we discuss the special case of the graphon Fourier transform for samples of an SBM. Namely, we present a method to construct the SBM Fourier basis using the model matrix and probability matrix associated with the SBM (Section 3.2), and describe how to implement this construction (Appendix A). We use this method in Section 4 to construct Fourier basis for SBM with a Cayley structure, and discuss how such Fourier bases can be transferred to samples of the SBM. In particular, we construct the SBM Fourier basis in the two cases of equal block sizes (Theorem 4.4) and arbitrary block sizes (Theorem 4.10). To establish the robustness of the Fourier transform from Theorem 4.4, we dedicate Section 4.2 to investigate the effect of the block size perturbations on the SBM Fourier transform. In Theorem 4.10, we apply these results to the case of a Cayley SBM with Abelian underlying group. It turns out that non-uniformity of block sizes can be interpreted as weighted Fourier transform (Remark 4.14). We conclude the paper with an example on ${\mathbb{Z}}_{5}$-SBMs in Section 5.
2 Notations and background
In this section, we collect the relevant information regarding the stochastic block model, graphons, and graph/graphon Fourier transforms.
[…]
2.1 Graphons and the stochastic block model
The stochastic block model (SBM) is a random graph model defined by a sequence of block sizes $k_{1},\ldots,k_{n}\in\mathbb{N}$ and a symmetric $n\times n$ probability matrix $A$. In this model, graphs are generated as follows: Initially, $N=\sum_{i=1}^{n}k_{i}$ vertices are created and partitioned into sets $B_{1},\dots,B_{n}$, referred to as blocks, according to the prescribed block sizes; then, for each $i,j\in[n]$, and for all $u\in B_{i}$, $v\in B_{j}$, edges $uv$ are independently added with probability $a_{i,j}$.
In order to create converging sequences of block model graphs, we will fix a probability measure $\mu$ on $[n]$, representing the fraction of vertices contained in each block. For a given $n\times n$ matrix $A$ and measure $\mu$ on $[n]$, we can then consider a sequence $\{G_{N}\}_{N}$ of graphs of increasing size formed according to the stochastic block model defined by probability matrix $A$ and block sizes $k_{i}=\mu_{i}N$ ($1\leq i\leq n$). To simplify the exposition, we assume that $\mu_{i}N$ is an integer for every $i$. We will say that $G_{N}\sim{\rm{SBM}}(A,\mu,N)$.
Remark 2.1.
Our exposition can easily be extended to remove the restriction that $\mu_{i}N$ is an integer. Namely, we can take $\{k_{i}\}_{i=1}^{n}$ so that $|k_{i}-\mu_{i}N|<1$ for all $i\in[n]$ and $\sum_{i=1}^{n}k_{i}=N$. The asymptotic results in this paper will remain true.
In [19], graphons were introduced as limit objects of convergent graph sequences. Graphons retain the large-scale structure of the graphs they represent, and provide a flexible framework for modeling large networks in a variety of applications (see e.g., [26, 9, 21, 14, 6]).
Definition 2.2.
A function $w:[0,1]^{2}\to[0,1]$ is called a graphon if $w$ is measurable and symmetric (i.e. $w(x,y)=w(y,x)$ almost everywhere).
The appropriate norm on the space of graphons is the cut norm, introduced in [10] and denoted by $\|\cdot\|_{\Box}$. For a measurable function $f:[0,1]^{2}\to{\mathbb{R}}$, the cut norm is defined as follows:
[…]
where $S$ and $T$ are taken over all measurable subsets of $[0,1]$.
Example 2.3 (Graphs and matrices as graphons).
An $N\times N$ matrix $A=[a_{i,j}]$ can be represented as a graphon $w_{A}$ by setting $w_{A}$ equal to $a_{i,j}$ on $[\frac{i-1}{N},\frac{i}{N})\times[\frac{j-1}{N},\frac{j}{N})$ for every $i,j\in[N]$. Any labeled graph can be represented by the ($\{0,1\}$-valued) graphon corresponding to its adjacency matrix.
Example 2.4 (SBMs as graphons).
An SBM given by $n\times n$ probability matrix $A$ and measure $\mu$ on $[n]$ can be presented as a graphon $w_{A,\mu}$ defined as follows. Let $\{I_{s}\}_{s=1}^{n}$ be a partition of $[0,1]$ into consecutive intervals so that, for all $s$, $|I_{s}|=\mu_{s}$. Then for all $1\leq s\leq t\leq n$ and for all $x\in I_{s}$, $y\in I_{t}$,
$w_{A,\mu}(x,y)=w_{A,\mu}(y,x)=a_{s,t}.$
Let $A$ and $\mu$ be as in Example 2.4. A sequence $\{G_{N}\}$, when $G_{N}\sim{{\rm{SBM}}}(A,\mu,N)$ and $N\to\infty$, almost surely forms a convergent graph sequence in the sense of graph limit theory (see [19]). Precisely, with suitable labelings of the graphs $\{G_{N}\}$, the sequence $\{w_{G_{N}}\}$ of associated graphons converges in cut norm to the graphon $w_{A,\mu}$.
Sometimes it will be useful to represent $w_{A,\mu}$ in the form of a matrix of a certain size. This motivates our introduction of the model matrix $W$.
Definition 2.5.
For $i\in[n]$, let $k_{i}=\mu_{i}N$. For any two positive integers $s,t$, let $J_{s,t}$ denote the $s\times t$ matrix consisting entirely of ones. Let $W=[a_{i,j}J_{k_{i},k_{j}}]$ be the matrix obtained by replacing every entry $a_{i,j}$ in $A$ by a block $a_{i,j}J_{k_{i},k_{j}}$. We call $W$ the model matrix of ${\rm{SBM}}(A,\mu,N)$. For any two vertices $u$ and $v$ in $G_{N}\sim{{\rm{SBM}}}(A,\mu,N)$, the probability that $u$ and $v$ are adjacent is given by $W_{u,v}$.
2.2 The graph/graphon Fourier transform
For a fixed graph $G$ with vertex set $V(G)$, a graph signal on $G$ is a function $f:V(G)\to{\mathbb{C}}$. If the vertex set $V(G)$ is labeled, say $\{v_{i}\}_{i=1}^{N}$, then the graph signal can be represented as a column vector $\left[f(v_{1}),f(v_{2}),\ldots,f(v_{N})\right]^{\rm T}$ in ${\mathbb{C}}^{N}$, where ${\rm T}$ denotes the matrix transpose.
To define the graph Fourier transform, we first assign a shift operator to the underlying graph. The transform is then given by the expansion of signals onto a fixed eigenbasis of this graph shift operator. Common choices for the shift operator include the graph adjacency matrix and the graph Laplacian. The adjacency matrix of a graph $G$ with $N$ nodes is a $0/1$-valued matrix $A_{G}$ of size $N$, whose $(i,j)$-th entry is 1 precisely when the vertices $v_{i}$ and $v_{j}$ are adjacent. The Laplacian of $G$, denoted by $L_{G}$, is the $N\times N$ matrix given by $L_{G}=D_{G}-A_{G}$, where $D_{G}$ is the diagonal matrix with entries $d_{ii}$ equal to the degree of vertex $v_{i}$. The selection of the graph shift operator significantly influences the properties of the resulting graph Fourier transform (see [27, Chapter 3]). For an overview of various shift operators used for developing graph Fourier transform, see [36] and references therein.
In this paper, we take the adjacency matrix as our graph shift operator. Our choice is due to the fact that the spectral features of the adjacency matrices associated with a converging sequence of graphs converge in an appropriate sense (see [37]). This phenomenon allows us to leverage the graph limit theory to produce consistent graph Fourier transforms. This is akin to the choice of shift operator in the definition of graphon Fourier transform proposed in [32] or the graphon neural networks in [33].
Being real symmetric matrices, adjacency matrices are unitarily diagonalizable. Let $A_{G}=U^{*}\Lambda U$, where $U$ is a unitary matrix and $\Gamma$ is the diagonal matrix whose diagonal entries are the eigenvalues of $A_{G}$. Given a graph signal $X\in{\mathbb{C}}^{N}$ on $G$, the graph Fourier transform of $X$ is defined as
(2.1) $\widehat{X}=UX.$
Given $\widehat{X}$, we can retrieve the original signal via the inverse Fourier transform defined as
(2.2) $X=U^{*}\widehat{X}.$
For a thorough discussion on graph Fourier transform and its applications, see for example [28, 27].
Similar to the graph Fourier transform, the graphon Fourier transform is derived from the spectral decomposition of the related graphon, or to be precise, the spectral decomposition of the associated operator. Every graphon $w:[0,1]^{2}\to[0,1]$ acts as the kernel of an integral operator on the Hilbert space $L^{2}[0,1]$ as follows:
[…]
An $L^{2}$ function is a $\lambda$-eigenfunction of $T_{w}$ (or $\lambda$-eigenfunction of $w$ for short) if $T_{w}f=\lambda f$ almost everywhere. For any graphon $w$, $T_{w}$ is a self-adjoint compact operator, and its operator norm is bounded by 1. Thus, $T_{w}$ has a countable spectrum lying in the interval $[-1,1]$ for which 0 is the only possible accumulation point. Let $n^{+}$ and $n^{-}$ in ${\mathbb{N}}_{*}:={\mathbb{N}}\cup\{\infty\}$ denote the number of positive and negative eigenvalues of $T_{w}$ respectively. We label the nonzero eigenvalues of $T_{w}$ as follows:
(2.3) $1\displaystyle 1\geq\lambda_{1}(w)\geq\lambda_{2}(w)\geq\ldots>0\ \mbox{ and }\ 0>\ldots\geq\lambda_{-2}(w)\geq\lambda_{-1}(w)\geq-$ 1
Note that if $n^{+}$ (resp. $n^{-}$) is finite, the corresponding chain of inequalities is a finite chain.
From the spectral theory for compact operators, we have that $L^{2}[0,1]$ admits an orthonormal basis containing eigenfunctions of $T_{w}$. Let $I_{w}\subseteq{\mathbb{Z}}\setminus\{0\}$ be the indices in (2.3) enumerating the nonzero (repeated) eigenvalues of $T_{w}$. Let $\{\phi_{i}\}_{i\in I_{w}}$ be an orthonormal collection of associated eigenfunctions. Then the spectral decomposition of $T_{w}$ is given as follows.
(2.4) $T_{w}=\sum_{i\in I_{w}}\lambda_{i}(w)\,\phi_{i}\otimes\phi_{i},$
where $\phi_{i}\otimes\phi_{i}$ denotes the rank-one projection defined as $(\phi_{i}\otimes\phi_{i})(\xi)=\langle\xi,\phi_{i}\rangle\phi_{i}$ for every $\xi\in L^{2}[0,1]$. Clearly, $\{\phi_{i}\}_{i\in I_{w}}$ forms an orthonormal basis for the image of $T_{w}$.
We can now generalize the definitions of graph signals and graph Fourier transform to the graphon setting. The concept of graphon Fourier transform was first introduced in [32] for graphons without any repeated eigenvalues. We quote the definition from a slightly generalized version proposed in [12], where the condition on eigenvalues was dropped.
Definition 2.6.
[12, Definition 3.4] Let $w:[0,1]^{2}\to[0,1]$ be a graphon, and consider the spectral decomposition of the associated integral operator as in (2.4).
  • (i)
    Any square-integrable measurable function $f\in L^{2}[0,1]$ is called a graphon signal on $w$.
  • (ii)
    The graphon Fourier transform of a signal $f$ is the projection of $f$ onto eigen-spaces of $T_{w}$. Precisely, for any distinct eigenvalue $\lambda$, let $I_{\lambda}$ denote the set of indices $i$ so that $\lambda_{i}=\lambda$. The Fourier projection of $f$ corresponding to $\lambda$, denoted by $\widehat{f}(\lambda)$, is defined as
    (2.5) $\widehat{f}(\lambda)=\sum_{i\in I_{\lambda}}(\phi_{i}\otimes\phi_{i})(f)=\sum_{i\in I_{\lambda}}\langle f,\phi_{i}\rangle\,\phi_{i}.$
    In addition, $\widehat{f}(0)$ is the projection of the signal onto the null space of $T_{w}$.
In this work, we use “Fourier projection” and “Fourier coefficient” interchangeably, with the understanding that the Fourier coefficient of a signal is a vector in general, rather than a scalar. In order to provide a common framework for signal processing on samples of a $w$-random graph, we need to view graph signals as elements of the space $L^{2}[0,1]$ of graphon signals.
Definition 2.7.
Let $Y$ be a signal on a graph $G$ with labeled vertex set $[N]$. We associate the step-function $f_{Y}\in L^{2}[0,1]$ to the graph signal $Y$ defined as:
(2.6) $f_{Y}(x)={\sqrt{N}}Y_{j}\text{ for all }\ x\in\left[\frac{j-1}{N},\frac{j}{N}\right],\,j\in[N].$
Here $\sqrt{N}$ is the scaling factor which ensures that $\|f_{Y}\|_{L^{2}[0,1]}=\|Y\|_{{\mathbb{C}}^{N}}$.
Remark 2.8.
The spectral features of an $N\times N$ matrix $A=[a_{i,j}]$ and its associated graphon $w_{A}$ (Example 2.3) are closely related. Let $X\in{\mathbb{C}}^{N}$ be a vector, and let $f_{X}$ denote the representation of $X$ as a function in $L^{2}[0,1]$ as given in (2.6). It is easy to observe the following:
[…]
Then we see that for any vector $Y\in{\mathbb{C}}^{N}$ and $\lambda\neq 0$, $f_{Y}$ is an eigenfunction of $T_{w_{A}}$ with eigenvalue $\lambda$ if and only if $Y$ is an eigenvector of $A$ with eigenvalue $\lambda N$. On the other hand, suppose $\lambda\neq 0$, and $f\in L^{2}[0,1]$ is a $\lambda$-eigenfunction of $T_{w_{A}}$. It is easy to see that $T_{w_{A}}f$ is constant on each interval $[\frac{j-1}{N},\frac{j}{N})$. Thus $f=\frac{1}{\lambda}T_{w_{A}}f$ is of the form of $f_{Y}$ for some $Y\in{\mathbb{C}}^{N}$. So if $\lambda\neq 0$, the map $Y\mapsto f_{Y}$ forms a one-to-one correspondence between $\lambda N$-eigenvectors of $A$ and $\lambda$-eigenfunctions of $T_{w_{A}}$.
Let $\{G_{n}\}_{n}$ be a sequence of labeled graphs of increasing size converging to the graphon $w$ in cut norm. For each $n$, let $\{\lambda_{i}{(G_{n})}\}_{i}$ be the sequence of eigenvalues of the adjacency matrix of $G_{n}$ ordered as in (2.3). By [18, Theorem 11.54], we have
$\lim_{n\to\infty}\frac{\lambda_{i}(G_{n})}{|V(G_{n})|}=\left\{\begin{array}[]{cc}\lambda_{i}(w)&\mbox{ if }i\in I_{w}\\ 0&\mbox{otherwise}\end{array}\right.$
We can now formulate the convergence of the GFT to the graphon Fourier transform, in the following sense. The following theorem was first proved in [32] for graphons that do not admit any repeated eigenvalues.
Theorem 2.9.
([12, Theorem 3.7]) Let $\{G_{n}\}$ be a graph sequence converging to a graphon $w$ in cut norm, and $\{Y_{n}\}$ be a sequence of graph signals on each of the $G_{n}$, such that the corresponding sequence of step-functions $\{f_{Y_{n}}\}_{n}$ converges to a graphon signal $f$ in $L^{2}[0,1]$. With the notation introduced in this section, we have that for each distinct nonzero eigenvalue $\lambda$ of $T_{w}$,
(2.7) $\left(\sum_{i\in I_{\lambda}}\phi^{n}_{i}\otimes\phi^{n}_{i}\right)(f)\rightarrow\widehat{f}(\lambda)\mbox{ in }L^{2}[0,1]\mbox{ as }n\rightarrow\infty,$
where each $\phi^{n}_{i}$ is an eigenvector of $G_{n}$ corresponding to $\lambda^{n}_{i}$, represented as a step-function.
If $\lambda=\lambda_{i}$ is an eigenvalue of $T_{w}$ of multiplicity 1, then this theorem implies that the sequence of $i$-th Fourier coefficients of signals $f_{n}$ on graphs $G_{n}$ converges, and the limit equals the (scalar) graphon Fourier coefficient $\widehat{f}(\lambda)$. For eigenvalues with higher multiplicity, this is not necessarily the case (see e.g. [12, Example 3.11]). Indeed, Theorem 2.9 implies that the projections onto eigenvectors corresponding to eigenvalues converging to the same value should be taken together. For example, if $\lambda$ has multiplicity 2 and $\lambda=\lambda_{1}=\lambda_{2}$, then the sequences $\left\{\frac{\lambda_{1}(G_{n})}{|V(G_{n})|}\right\}_{n}$ and $\left\{\frac{\lambda_{2}(G_{n})}{|V(G_{n})|}\right\}_{n}$ both converge to $\lambda$. Theorem 2.9 then states that the projection of the signal $f_{n}$ onto the space spanned by the eigenvectors of $G_{n}$ corresponding to $\lambda_{1}(G_{n})$ and $\lambda_{2}(G_{n})$ converges to the projection of $f$ onto the eigenspace of $\lambda$.
3 The graphon Fourier transform for the SBM
In this section, we particularize the concept of the graphon Fourier transform to the case of stochastic block models taking the associated model matrix into account. This prepares the ground for our discussion for stochastic block models with Cayley structure.
Consider a sequence $\{G_{N}\}$ of graphs of increasing size so that each $G_{N}$ is sampled from a stochastic block model with probability matrix $A$ and measure $\mu$. As seen earlier, $\{G_{N}\}$ converges almost surely to the graphon $w_{A,\mu}$ associated with the SBM. The graphon approach to GSP is to use the graphon basis for signal processing on all graphs in the sequence, instead of using the particular eigenbasis of each individual graph. In other words, the SBM-driven Fourier transform is based on the eigenbasis of $w_{A,\mu}$. In this section, we discuss how to obtain this basis and analyze its sensitivity to changes in the block sizes.
3.1 SBM and model matrix
Let $G$ be a graph sampled from a stochastic block model ${\rm{SBM}}(A,\mu,N)$. Signals on $G$ are vectors in ${\mathbb{C}}^{N}$. To apply the SBM-driven Fourier transform directly to a signal $X$, we use an eigenbasis of the model matrix $W$ (Definition 2.5). Under our assumption that $\mu_{i}N\in{\mathbb{N}}$ for all $i\in[n]$, we have that $w_{A,\mu}$ is the graphon representation of $W$ (see Example 2.3). Let $X\in{\mathbb{C}}^{N}$ be a signal on $G$ and $\lambda$ a non-zero eigenvalue of $T_{w_{A,\mu}}$. The graphon Fourier transform of $f_{X}$ as given in Definition 2.6 is the projection of $f_{X}$ onto the $\lambda$-eigenspace of $T_{w_{A,\mu}}$. Using the correspondence between eigenspaces of $W$ and $T_{w_{A,\mu}}$ (Remark 2.8), we can express this projection in terms of vectors in ${\mathbb{C}}^{N}$ as $\widehat{X}(\lambda)=P_{\lambda}(X)$, where $P_{\lambda}$ is the projection, in ${\mathbb{C}}^{N}$, onto the $\lambda N$-eigenspace of $W$.
The Fourier coefficient for eigenvalue zero has special significance. In graph signal processing, the adjacency matrix plays the role of the shift operator, which models the spread of a signal along the edges of the graph. The eigenvectors corresponding to large eigenvalues (in magnitude) are considered smooth with respect to the shift. In contrast, eigenvectors of eigenvalue zero are highly non-smooth. Therefore, the Fourier projection corresponding to eigenvalue zero may be viewed as the noise present in the signal.
In the graphon Fourier approach, the Fourier projection of $X$ corresponding to eigenvalue 0 is the projection of $f_{X}$ onto the kernel of $T_{w_{A,\mu}}$. However, contrary to other eigenspaces, the kernel of $T_{w_{A,\mu}}$ and of $W$ do not correspond exactly. It follows easily from the definition of $w_{A,\mu}$ (and also from Proposition 3.4 below) that $T_{w_{A,\mu}}$ has rank at most $n$. Therefore, $T_{w_{A,\mu}}$ has eigenvalue zero with infinite multiplicity. The rank of $W$ equals that of $T_{w_{A,\mu}}$, so the dimension of the kernel of $W$ is finite but grows with $N$, more precisely, we have $N-n\leq\dim(\ker{W})\leq N$.
Instead of having to compute an eigenbasis for 0, we can use the orthogonal complement of the 0-eigenspace to define the projection $P_{0}(X)$ indirectly as
(3.1) $P_{0}(X)=X-\sum_{\lambda\not=0}P_{\lambda}(X),$
where the sum ranges over all non-zero eigenvalues of $W$. Since $w_{A,\mu}$ has a finite number of non-zero eigenvalues, the convergence result of Theorem 2.9 holds for $P_{0}$ as well.
3.2 Computing the SBM Fourier basis
In this section we will see that the eigen-decomposition of the model matrix $W$ is closely related to that of a much smaller matrix $A_{\mu}$, which we now define.
Definition 3.1.
Let ${\rm{SBM}}(A,\mu,N)$ be an SBM with an $n\times n$ probability matrix $A$. We introduce the following matrices.
  1. (a)
    The diagonal matrix $M:={\rm diag}(\mu_{1},\ldots,\mu_{n})$ is called the weight matrix.
  2. (b)
    The $n\times n$ matrix $A_{\mu}:=\sqrt{M}A\sqrt{M}$ is called the weighted probability matrix.
  3. (c)
    The $N\times n$ matrix $D$ is defined in block form as $D={\rm diag}(J_{k_{1},1},J_{k_{2},1},\ldots,J_{k_{n},1})$, where $J_{k_{i},1}$ is the column vector in ${\mathbb{C}}^{k_{i}}$ whose entries are all ones.
  4. (d)
    The $N\times n$ matrix $V$ is defined as $V:=\frac{1}{\sqrt{N}}DM^{-\frac{1}{2}}.$
Note that for any vector $X\in{\mathbb{C}}^{n}$, $DX$ is the “blow-up”vector in the high-dimensional space ${\mathbb{C}}^{N}$ obtained from $X$ by repeating each entry $X_{i}$ exactly $k_{i}$ times.
Lemma 3.2.
With notations as given above in Definition 3.1, and model matrix $W$ as in Definition 2.5, we have the following.
  1. (a)
    $W=DAD^{\rm{T}}$ and $\frac{1}{N}D^{\rm{T}}D=M$.
  2. (b)
    The operator $V$ is an isometry, i.e., $V^{\rm{T}}V=I$.
  3. (c)
    $VA_{\mu}V^{\rm{T}}=\frac{1}{N}W$ and $A_{\mu}=\frac{1}{N}(V^{\rm{T}}WV)$.
  4. (d)
    $\|D\|_{\rm{opr}}=\sqrt{N\max\{\mu_{1},\ldots,\mu_{n}\}}$.
Proof 3.3.
[…]
$\langle VX,VY\rangle_{{\mathbb{C}}^{N}}$ $=$ $\langle\frac{1}{\sqrt{N}}DM^{-1/2}X,\frac{1}{\sqrt{N}}DM^{-1/2}Y\rangle_{{\mathbb{C}}^{N}}=\langle\frac{1}{N}D^{\rm{T}}DM^{-1/2}X,M^{-1/2}Y\rangle_{{\mathbb{C}}^{n}}$
$=$ $\langle M^{-1/2}MM^{-1/2}X,Y\rangle_{{\mathbb{C}}^{n}}=\langle X,Y\rangle_{{\mathbb{C}}^{n}}.$
[…]
$\|D\|^{2}_{\rm{opr}}=\|D^{\rm{T}}D\|_{\rm{opr}}=N\|M\|_{\rm{opr}}=N\max_{1\leq i\leq n}\mu_{i}.$
It follows from Lemma 3.2 that $W$ is a multiple of the conjugation of $A_{\mu}$ by an isometry. We will now see that the spectral behavior of $A_{\mu}$ and $W$ are closely related.
Proposition 3.4.
Let $W$, $M$ and $A_{\mu}$ be the model matrix, the weight matrix, and the weighted probability matrix of an SBM defined by $(A,\mu,N)$, and let $D$ and $V$ be as given in Definition 3.1. Let $\lambda\in{\mathbb{R}}$, and $X\in{\mathbb{C}}^{n}$ and $Y\in{\mathbb{C}}^{N}$ be unit vectors. Then,
  1. (a)
    Suppose $\lambda\neq 0$. If $Y$ is a unit $\lambda$-eigenvector of $W$ then $Y\in{\rm range}(V)$, and $X=V^{\rm{T}}Y$ is a unit $\frac{\lambda}{N}$-eigenvector of $A_{\mu}$.
  2. (b)
    If $X$ is a unit $\frac{\lambda}{N}$-eigenvector of $A_{\mu}$ then $Y=VX$ is a unit $\lambda$-eigenvector of $W$.
  3. (c)
    $\lambda$ is a nonzero eigenvalue of $W$ of multiplicity $t$ if and only if $\frac{\lambda}{N}$ is a nonzero eigenvalue of $A_{\mu}$ of multiplicity $t$.
Proof 3.5.
To prove Item (a), suppose that $Y$ is a unit $\lambda$-eigenvector of $W$, i.e, $WY=\lambda Y$. Since $W=N(VA_{\mu}V^{\rm{T}})$ (by Lemma 3.2 Item (c)) and $Y=\frac{1}{\lambda}WY$, we get $Y\in{\rm range}(V)$. Consequently, since $V^{\rm{T}}$ restricted to ${\rm range}(V)$ is an isometry, the vector $X=V^{\rm{T}}Y$ is a unit vector as well. More precisely, since $Y=V\tilde{Y}$ for some $\tilde{Y}\in{\mathbb{C}}^{n}$ and $V$ is an isometry, we have
$\|X\|=\|V^{\rm{T}}Y\|=\|V^{\rm{T}}V\tilde{Y}\|=\|\tilde{Y}\|=\|V\tilde{Y}\|=\|Y\|=1.$
Using Lemma 3.2 Item (c), we have
$\frac{1}{N}V^{\rm{T}}WY=\frac{1}{N}{V^{\rm{T}}}(NVA_{\mu}V^{\rm{T}})Y=A_{\mu}V^{\rm{T}}Y=A_{\mu}X.$
Since $Y$ is a $\lambda$-eigenvector of $W$, we have $\frac{1}{N}V^{\rm{T}}WY=\frac{\lambda}{N}V^{\rm{T}}Y=\frac{\lambda}{N}X$. Putting these two identities together, we get that $X$ is a unit $\frac{\lambda}{N}$-eigenvector of $A_{\mu}$.
To prove Item (b), assume that $X$ is a unit vector satisfying $A_{\mu}X=\frac{\lambda}{N}X$. Since $V$ is an isometry, $Y=VX$ is a unit vector as well. Moreover, using Lemma 3.2 Item (c) again, we have
$WY=(NVA_{\mu}V^{\rm{T}})(VX)=NVA_{\mu}X=NV(\frac{\lambda}{N}X)=\lambda Y.$
To prove Item (c), suppose $\lambda\neq 0$ is an eigenvalue of $W$, and $\{Y_{1},\ldots,Y_{t}\}$ is an orthonormal basis for the associated $\lambda$-eigenspace. By Item (a), the set $\{V^{\rm{T}}Y_{1},\ldots,V^{\rm{T}}Y_{t}\}$ is a subset of the $\frac{\lambda}{N}$-eigenspace of $A_{\mu}$. Moreover, since each $Y_{i}$ belongs to the range of $V$ (by Item (a)), and using the fact that the operator $V^{\rm{T}}$ when restricted to the subspace ${\rm range}(V)$ is an isometry, we observe that the set $\{V^{\rm{T}}Y_{1},\ldots,V^{\rm{T}}Y_{t}\}$ is orthonormal as well. Thus the multiplicity of $\frac{\lambda}{N}$ as an eigenvalue of $A_{\mu}$ is at least $t$. Conversely, let $\frac{\lambda}{N}$ be a nonzero eigenvalue of $A_{\mu}$ with orthonormal eigenbasis $\{X_{1},\ldots,X_{s}\}$. By Item (b), the set $\{VX_{1},\ldots,VX_{s}\}$ is an orthogonal family of $\lambda$-eigenvectors of $W$. Therefore the multiplicity of $\frac{\lambda}{N}$ as an eigenvalue of $A_{\mu}$, is at most $t$.
With this proposition in hand, we can reduce computing the eigen-decomposition of the range of the matrix $W$ to the eigen-decomposition of the smaller matrix $A_{\mu}$. Note that $W$ is $N\times N$ (and $N$ usually tends to infinity), whereas the size of $A_{\mu}$ is equal to the number of the blocks in SBM. This correspondence extends to the nonzero eigenvalues/eigenvectors of the graphon operator $T_{w_{A,\mu}}$ as well, since $\lambda\neq 0$ is an eigenvalue of $T_{w_{A,\mu}}$ if and only if $\lambda N$ is an eigenvalue of $W$ (Remark 2.8). The following corollary then follows directly from Proposition 3.4.
Corollary 3.6.
The nonzero eigenvalues of the graphon $w_{A,\mu}$ are exactly the nonzero eigenvalues of $A_{\mu}$.
The theory developed in this section leads to a straightforward algorithm for computing SBM Fourier basis. The details of this algorithm can be found in Appendix A.
4 Fourier basis for SBM with group symmetries
In this section, we discuss stochastic block models whose structure is informed by certain group symmetries. We formalize the group theoretic symmetries of an SBM through its probability matrix as follows.
Definition 4.1.
An $n\times n$ matrix $A$ is called a Cayley matrix on a group ${\mathbb{G}}$ if $|{\mathbb{G}}|=n$ and there exists a connection function $f:{\mathbb{G}}\rightarrow[0,1]$ so that $a_{i,j}=f(g_{i}^{-1}g_{j})$ for all $i,j\in[n]$. The function $f$ must be inverse-invariant, i.e. $f(x)=f(x^{-1})$. An ${\rm{SBM}}(A,\mu,n)$ is called a ${\mathbb{G}}$-SBM, if $A$ is a Cayley matrix on the group ${\mathbb{G}}$.
A symmetric matrix can be viewed as a weighted graph, where the edge between vertices $i$ and $j$ has weight $a_{ij}$. With this interpretation, the definition of a Cayley matrix coincides with that of a Cayley graph. This fact has inspired our choice of terminology in Definition 4.1.
4.1 ${\mathbb{G}}$-SBM with uniform measure
Let ${\rm{SBM}}(A,\mu,N)$ be a ${\mathbb{G}}$-SBM with the connection function $f:{\mathbb{G}}\to[0,1]$. Here, we consider the special case where $\mu$ is the uniform measure on $[n]$. Under this condition, $A_{\mu}=\frac{1}{n}A$, which is a Cayley matrix, and the associated graphon $w_{A,\mu}$ becomes a Cayley graphon. Cayley graphons are generalizations of Cayley graphs, and were first introduced in [20]. The Cayley graphons of finite groups are precisely the graphons associated with a ${\mathbb{G}}$-SBM with uniform measure. For such SBM, the graphon Fourier basis can be derived directly from the Cayley matrix $A$. Eigen-decompositions of Cayley matrices are well-understood when the group is Abelian ([2]) or quasi-Abelian ([30]). Representation theory of groups has been used to construct ‘preferred’ Fourier bases for signal processing on Cayley graphs and graphons (see e.g. [11, 12, 4, 7]).
Using harmonic analysis, it is easy to find a graphon Fourier basis for a ${\mathbb{G}}$-SBM where ${\mathbb{G}}$ is Abelian. Let ${\mathbb{G}}=\{g_{1},\ldots,g_{n}\}$ be an Abelian group, and ${\mathbb{T}}$ be the multiplicative group of complex numbers with modulus 1 represented as ${\mathbb{T}}=\{e^{2\pi ix}:0\leq x<1\}$. Characters of ${\mathbb{G}}$ are maps $\chi:{\mathbb{G}}\to{\mathbb{T}}$ satisfying $\chi(gh)=\chi(g)\chi(h)$ for all $g,h\in{\mathbb{G}}$. We denote the collection of all characters of ${\mathbb{G}}$ by $\widehat{{\mathbb{G}}}$. The set of normalized characters $\left\{\frac{\chi}{\sqrt{n}}:\ \chi\in\widehat{{\mathbb{G}}}\right\}$ forms an orthonormal basis for the vector space ${\mathbb{C}}^{n}$. Here, we think of ${\mathbb{C}}^{n}$ as vectors indexed by elements of ${\mathbb{G}}$. This allows us to identify functions on ${\mathbb{G}}$ with vectors in ${\mathbb{C}}^{n}$.
The characters of a group give a diagonalization of a Cayley matrix on ${\mathbb{G}}$ as follows. For a proof, see e.g., [2, 4].
Lemma 4.2 (Diagonalization of Cayley matrices).
Let $\mathbb{G}$ be an Abelian group, and $f:{\mathbb{G}}\rightarrow[0,1]$ be a connection function. Let $A$ be the Cayley matrix on ${\mathbb{G}}$ defined by $f$. Let $\widehat{{\mathbb{G}}}=\{\chi_{1},\ldots,\chi_{n}\}$ be the set of characters of ${\mathbb{G}}$. Then every character $\chi_{i}\in\widehat{{\mathbb{G}}}$ is an eigenvector of $A$ associated with eigenvalue $\lambda_{i}=\Sigma_{x\in{\mathbb{G}}}f(x)\overline{\chi_{i}(x)}$. Consequently, the unitary matrix $U$ defined as $U:=\left[\frac{\chi_{j}(g_{i})}{\sqrt{n}}\right]_{i,j}$ diagonalizes $A$:
$U^{*}AU=\Gamma:={\rm diag}\left({\sum\limits_{x\in{\mathbb{G}}}f(x)\overline{\chi_{1}(x)}},\ldots,\sum\limits_{x\in{\mathbb{G}}}f(x)\overline{\chi_{n}(x)}\right).$
The characters of ${\mathbb{G}}$ thus provide an eigenbasis for a Cayley matrix on a group ${\mathbb{G}}$ no matter which connection function $f$ is used.
Remark 4.3.
For $\chi\in\widehat{{\mathbb{G}}}$, let $\overline{\chi}$ denote the character defined by $\overline{\chi}(x)=\overline{\chi(x)}$ for all $x\in{\mathbb{G}}$. Since $\overline{\chi(x)}=\chi(x^{-1})$ and $f$ is inverse-invariant, we have that the eigenvalues associated with $\chi$ and $\overline{\chi}$ are equal. Therefore, $A$ has a repeated eigenvalue for each character $\chi$ so that $\chi\not=\overline{\chi}$. Moreover, $\chi+\overline{\chi}$ and $i(\chi-\overline{\chi})$ provide a real-valued pair of eigenvectors that span the same space as $\chi$ and $\overline{\chi}$.
Theorem 4.4 (GFT for samples of ${\mathbb{G}}$-SBM with uniform measure).
Let $W$ be the model matrix of ${\rm{SBM}}(A,\mu,N)$, where $A$ is a Cayley matrix on an Abelian group ${\mathbb{G}}$ with the connection function $f$. Suppose $\mu$ is the uniform measure on $[n]$. For unit vector $Y\in{\mathbb{C}}^{N}$ and $\lambda\neq 0$,
$Y\mbox{ is a }\lambda\mbox{-eigenvector of }W\iff\ \exists\chi\in\widehat{G}\ \mbox{ s.t. }\ \frac{\lambda}{N}=\frac{1}{n}\sum\limits_{x\in{\mathbb{G}}}f(x)\overline{\chi(x)}\ \ \&\ \ Y=\frac{1}{\sqrt{N}}D\chi,$
where $D$ is given in Definition 3.1.
Proof 4.5.
Since $\mu$ is a uniform measure, we have $A_{\mu}=\frac{1}{n}A$, $M=\frac{1}{n}I$, and $V=\sqrt{\frac{n}{N}}D$. By Lemma 4.2, the collection $\{\frac{1}{n}\sum\limits_{x\in{\mathbb{G}}}f(x)\overline{\chi(x)}\}_{\chi\in\widehat{{\mathbb{G}}}}$ includes all eigenvalues of $A_{\mu}$ with corresponding orthonormal eigenbasis $\{\frac{1}{\sqrt{n}}\chi\}_{\chi\in\widehat{{\mathbb{G}}}}$. The statement then follows directly from the relation between the eigen-decompositions of $A_{\mu}$ and $W$ given in Proposition 3.4.
In this paper, we restrict our attention to Cayley graphons on Abelian groups. If the underlying group is not Abelian, the representation theory of the group may be used to develop a specific basis for the graphon Fourier transform. This basis can then be used to provide an instance-independent framework for graph signal processing on the samples of the SBM. Details can be found in [12]; for applications on the symmetric group, see [3, 7].
4.2 General robustness results under block size perturbation
In this section, we explore how the graphon Fourier transform defined on samples of an SBM varies when the block ratios in the SBM are adjusted slightly. We remark that the results in this subsection are valid for general SBMs. We apply these general results to the Cayley setting in the subsequent sections.
As shown in Section 3.2, the eigen-decomposition of the $n\times n$ weighted probability matrix $A_{\mu}$ leads to a basis for the graphon Fourier transform for the SBM defined by $A$ and $\mu$. This point of view allows us to use tools from matrix perturbation theory to control the effect of small variations in block sizes on the resulting graphon Fourier transform.
Let $A$ be a probability matrix of size $n$. Fix $N\in{\mathbb{N}}$ and $\epsilon>0$. Let $\mu,\mu^{\prime}$ be probability measures on $[n]$ so that $\mu^{\prime}_{i}=\mu_{i}(1+\epsilon_{i})$ and $|\epsilon_{i}|\leq\epsilon$ for all $i\in[n]$. Let $W,W^{\prime}$ be the model matrices of ${\rm{SBM}}(A,\mu,N)$ and ${\rm{SBM}}(A,\mu^{\prime},N)$, respectively. Let $\{Y_{i}\}_{i\in I_{A,\mu}}$ and $\{Y_{i}^{\prime}\}_{i\in I_{A,\mu^{\prime}}}$ be orthonormal sets of eigenvectors associated with nonzero eigenvalues of $W$ and $W^{\prime}$ ordered as in (2.3), respectively.
The following theorem presents an upper bound on the change to $\widehat{X}$ if we use the basis $\{Y^{\prime}_{i}\}$ derived from the slightly different measure $\mu^{\prime}$, instead of the true basis $\{Y_{i}\}$. First, for any eigenvalue $\lambda$ of $A_{\mu}$ we define $\gamma(\lambda)$ as the gap between $\lambda$ and the other eigenvalues, i.e.,
$\gamma(\lambda)=\min\{|\lambda-\lambda_{i}|:\lambda_{i}\not=\lambda\}.$
Theorem 4.6.
Consider any nonzero eigenvalue $\lambda$ of $A_{\mu}$, with multiplicity $d$. Then for any signal $X\in{\mathbb{C}}^{N}$,
(4.1) $\|\widehat{X}(\lambda)-\sum_{i\in I_{\lambda}}\langle X,Y^{\prime}_{i}\rangle_{{{\mathbb{C}}}^{N}}Y^{\prime}_{i}\|_{{\mathbb{C}}^{N}}\leq\left(\frac{2^{5/2}d^{1/2}\|A_{\mu}\|_{{\rm{opr}}}}{\gamma(\lambda)}n\epsilon+\frac{2\sqrt{3}}{\sqrt{\mu_{\min}}}\sqrt{n\epsilon}\right)\|X\|_{{\mathbb{C}}^{N}}.$
The proof of Theorem 4.6 depends heavily on the Davis-Kahan theorem on matrix perturbations. The relevant background can be found in Appendix B. First, we prove a necessary lemma that shows how the isometry $V$ is affected by the change in measure.
Lemma 4.7.
Let $V:{\mathbb{C}}^{n}\rightarrow{\mathbb{C}}^{N}$ (resp. $V^{\prime}:{\mathbb{C}}^{n}\rightarrow{\mathbb{C}}^{N}$) be the isometry given in Definition 3.1 for $\mu$ (resp. ${\mu^{\prime}}$). Then, letting $\mu_{\min}:=\min\{\mu_{1},\ldots,\mu_{n}\}$, we have
$\|V-V^{\prime}\|_{\rm{opr}}\leq\sqrt{\frac{3n}{\mu_{\min}}}\sqrt{\epsilon}.$
Proof 4.8.
Let $D,M$ (resp. $D^{\prime},M^{\prime}$) be the matrices associated with $\mu$ (resp. $\mu^{\prime}$) as in Definition 3.1. Recall that $k_{i}=\mu_{i}N$ and $k_{i}^{\prime}=\mu^{\prime}_{i}N$. Using the structure of $D$ and $D^{\prime}$, it is easy to observe that the $j$’th column of $D-D^{\prime}$ has nonzero elements in at most $|\sum_{i=1}^{j-1}k_{i}-\sum_{i=1}^{j-1}k^{\prime}_{i}|+|\sum_{i=1}^{j}k_{i}-\sum_{i=1}^{j}k^{\prime}_{i}|$ entries. Given that the matrix $D-D^{\prime}$ has $n$ columns, and its nonzero entries are either $1$ or $-1$, we have
(4.2) $\|D-D^{\prime}\|_{\rm{opr}}$$\leq$ $\|D-D^{\prime}\|_{F}\leq\sqrt{n\left(\left|\sum_{i=1}^{j-1}k_{i}-\sum_{i=1}^{j-1}k^{\prime}_{i}\right|+\left|\sum_{i=1}^{j}k_{i}-\sum_{i=1}^{j}k^{\prime}_{i}\right|\right)}$
$\leq$ $\sqrt{2n\sum_{i=1}^{n}|k_{i}-k^{\prime}_{i}|}=\sqrt{2n\sum_{i=1}^{n}|\epsilon_{i}|\mu_{i}N}\leq\sqrt{2n\epsilon N\sum_{i=1}^{n}\mu_{i}}=\sqrt{2\epsilon nN},$
where in the last equality we used the fact that $\sum_{i=1}^{n}\mu_{i}=1$. Next, note that $\|M^{-\frac{1}{2}}\|_{\rm{opr}}=\frac{1}{\sqrt{\mu_{\min}}}$ and
$\|M^{-\frac{1}{2}}-M^{\prime-\frac{1}{2}}\|_{\rm{opr}}$ $=$ $\max\left\{\left|\frac{1}{\sqrt{\mu_{i}}}-\frac{1}{\sqrt{\mu_{i}(1+\epsilon_{i})}}\right|:i\in[n]\right\}$
$=$ $\max\left\{\frac{1}{\sqrt{\mu_{i}}}\left|\frac{1-\sqrt{1+\epsilon_{i}}}{\sqrt{1+\epsilon_{i}}}\right|:i\in[n]\right\}\leq\frac{\epsilon}{2\sqrt{\mu_{\min}}}.$
These facts, together with (4.2) and Lemma 3.2 Item (d), lead to the desired estimates:
$\|V-V^{\prime}\|_{\rm{opr}}$ $=$ $\frac{1}{\sqrt{N}}\|DM^{-\frac{1}{2}}-D^{\prime}M^{\prime-\frac{1}{2}}\|_{\rm{opr}}$
$\leq$ $\frac{1}{\sqrt{N}}\left(\|D-D^{\prime}\|_{\rm{opr}}\|M^{-\frac{1}{2}}\|_{\rm{opr}}+\|D^{\prime}\|_{\rm{opr}}\|M^{-\frac{1}{2}}-M^{\prime-\frac{1}{2}}\|_{\rm{opr}}\right)$
$\leq$ $\frac{1}{\sqrt{N}}\left(\sqrt{2\epsilon nN}\frac{1}{\sqrt{\mu_{\min}}}+\sqrt{N}\frac{\epsilon}{2\sqrt{\mu_{\min}}}\right)\leq\frac{\sqrt{3n}}{\sqrt{\mu_{\min}}}\sqrt{\epsilon}.$
Proof 4.9 (Proof of Theorem 4.6).
Since $\mu^{\prime}$ is a probability measure on $[n]$, we have that $\sum_{i=1}^{n}\epsilon_{i}=0$. Note that
$A_{\mu^{\prime}}-A_{\mu}=\left[\sqrt{\mu_{i}}\sqrt{\mu_{j}}{a_{ij}}(\sqrt{1+\epsilon_{i}}\sqrt{1+\epsilon_{j}}-1)\right]_{i,j}=A_{\mu}\circ E,$
where $E=\left[\sqrt{1+\epsilon_{i}}\sqrt{1+\epsilon_{j}}-1\right]_{i,j}$ and $\circ$ denotes the Schur (or Hadamard) product of matrices. Let $Z,{\mathbf{1}}\in{\mathbb{C}}^{n}$ be defined as $Z=[\sqrt{1+\epsilon_{1}},\ldots,\sqrt{1+\epsilon_{n}}]^{\rm{T}}$, ${\mathbf{1}}=[1,\ldots,1]^{\rm{T}}$, and note that $E=(Z-\mathbf{1})Z^{\rm{T}}+\mathbf{1}(Z^{\rm{T}}-\mathbf{1}^{\rm{T}})$. Moreover, observe that $\|Z\|_{{\mathbb{C}}^{n}}\leq\sqrt{n(1+\epsilon)}$ and $\|Z-\mathbf{1}\|_{{\mathbb{C}}^{n}}\leq\sqrt{n}(\sqrt{1+\epsilon}-1)\leq\sqrt{n}(\frac{\epsilon}{2})$. Since the operator norm is submultiplicative with respect to the Schur product of matrices, using the above norm estimates we get
$\|A_{\mu^{\prime}}-A_{\mu}\|_{\rm{opr}}$$\leq$ $\|A_{\mu}\|_{\rm{opr}}\|E\|_{\rm{opr}}$
$\leq$ $\|A_{\mu}\|_{\rm{opr}}(\|Z-\mathbf{1}\|_{{\mathbb{C}}^{n}}\|Z\|_{{\mathbb{C}}^{n}}+\|\mathbf{1}\|_{{\mathbb{C}}^{n}}\|Z-\mathbf{1}\|_{{\mathbb{C}}^{n}})$
$\leq$ $n(\frac{\epsilon}{2})(\sqrt{1+\epsilon}+1)\|A_{\mu}\|_{\rm{opr}}.$
Since $\epsilon\leq 1$, we have the upper bound $\|A_{\mu^{\prime}}-A_{\mu}\|_{\rm{opr}}\leq 2\epsilon n\|A_{\mu}\|_{\rm{opr}}$.
Let $V:{\mathbb{C}}^{n}\rightarrow{\mathbb{C}}^{N}$ (resp. $V^{\prime}:{\mathbb{C}}^{n}\rightarrow{\mathbb{C}}^{N}$) be the isometry given in Definition 3.1 for $\mu$ (resp. ${\mu^{\prime}}$), and set $X_{i}=V^{\rm{T}}Y_{i}$ and $X^{\prime}_{i}={V^{\prime}}^{\rm{T}}Y^{\prime}_{i}$. By Proposition 3.4, $\{X_{i}\}_{i\in I_{A,\mu}}$ (resp. $\{X^{\prime}_{i}\}_{i\in I_{A,\mu^{\prime}}}$) is an orthonormal set of eigenvectors of $A_{\mu}$ (resp. $A_{\mu^{\prime}}$) associated with nonzero eigenvalues, indexed as in (2.3). Since all matrices $A_{\mu}$, $W$, $A_{\mu^{\prime}}$ and $W^{\prime}$ are symmetric and real-valued, their corresponding eigenbases can be chosen to be real-valued vectors. Fix a nonzero eigenvalue $\lambda$ of $A_{\mu}$ of multiplicity $d$ so that $\lambda=\lambda_{r}=\dots=\lambda_{s}$, where $d=s-r+1$. Let ${\mathcal{E}}={\rm span}\{X_{r},\dots,X_{s}\}$ and ${\mathcal{E}}^{\prime}={\rm span}\{X^{\prime}_{r},\dots,X^{\prime}_{s}\}$. Let $P_{\mathcal{E}}$ and $P_{{\mathcal{E}}^{\prime}}$ be the orthogonal projections, in ${\mathbb{R}}^{n}$, on ${\mathcal{E}}$ and ${\mathcal{E}}^{\prime}$, respectively. Then, by Corollary B.2, we have that
(4.3) $\|P_{\mathcal{E}}-P_{{{\mathcal{E}}}^{\prime}}\|_{{\rm{opr}}}\leq\|P_{\mathcal{E}}-P_{{{\mathcal{E}}}^{\prime}}\|_{F}\leq\frac{2^{3/2}d^{1/2}\|A_{\mu}-A_{\mu^{\prime}}\|_{{\rm{opr}}}}{\min\{\lambda_{r-1}-\lambda,\lambda-\lambda_{s+1}\}}\leq\frac{2^{5/2}d^{1/2}\epsilon n\|A_{\mu}\|_{\rm{opr}}}{\gamma(\lambda)}.$
Note that the projection $P_{\mathcal{E}}:{\mathbb{C}}^{n}\rightarrow{\mathbb{C}}^{n}$ is defined as
$P_{\mathcal{E}}(Z)=\sum_{i=r}^{s}\langle Z,X_{i}\rangle X_{i}=\sum_{i=r}^{s}(X_{i}^{*}Z)X_{i}=\sum_{i=r}^{s}X_{i}(X_{i}^{*}Z)=(\sum_{i=r}^{s}X_{i}X_{i}^{*})Z.$
So, when represented in matrix form, we have that $P_{\mathcal{E}}=\sum_{i=r}^{s}X_{i}X_{i}^{*}$; a similar formula holds for $P_{{\mathcal{E}}^{\prime}}$. Let $P_{E}$ and $P_{E^{\prime}}$ be the orthogonal projections on $E={\rm span}\{Y_{r},\dots,Y_{s}\}$ and $E^{\prime}={\rm span}\{Y^{\prime}_{r},\dots,Y^{\prime}_{s}\}$ respectively. By Proposition 3.4 (a), each $Y_{i}$ (resp. $Y_{i}^{\prime}$) is in the image of $V$ (resp. $V^{\prime}$), so $VX_{i}=VV^{\rm{T}}Y_{i}=Y_{i}$ and similarly $V^{\prime}X^{\prime}_{i}=Y^{\prime}_{i}$. Expressing $P_{E}$ in matrix form, we have that
$P_{E}=\sum_{i=r}^{s}(VX_{i})(VX_{i})^{*}=\sum_{i=r}^{s}VX_{i}\,X_{i}^{*}V^{*}=V\left(\sum_{i=r}^{s}X_{i}\,X_{i}^{*}\right)V^{*}=VP_{\mathcal{E}}V^{*}.$
Applying the same argument, we obtain that $P_{E^{\prime}}=V^{\prime}P_{{\mathcal{E}}^{\prime}}{V^{\prime}}^{*}$. So
$P_{E}-P_{E^{\prime}}=VP_{\mathcal{E}}V^{*}-V^{\prime}P_{{\mathcal{E}}^{\prime}}{V^{\prime}}^{*}=V(P_{\mathcal{E}}-P_{{\mathcal{E}}^{\prime}})V^{*}+(V-V^{\prime})P_{{\mathcal{E}}^{\prime}}V^{*}+V^{\prime}P_{{\mathcal{E}}^{\prime}}(V^{*}-{V^{\prime}}^{*}),$
which implies that
$\|P_{E}-P_{E^{\prime}}\|_{{\rm{opr}}}\leq\|P_{\mathcal{E}}-P_{{\mathcal{E}}^{\prime}}\|_{{\rm{opr}}}+\|V-V^{\prime}\|_{{\rm{opr}}}\|P_{{\mathcal{E}}^{\prime}}\|_{{\rm{opr}}}+\|P_{{\mathcal{E}}^{\prime}}\|_{{\rm{opr}}}\|V^{*}-{V^{\prime}}^{*}\|_{{\rm{opr}}},$
where we used the facts that operator norm is submultiplicative, and $\|V\|_{{\rm{opr}}}=\|V^{*}\|_{{\rm{opr}}}=1$, as $V$ is an isometry. Furthermore, since any projection is contractive, we get
(4.4) $\|P_{E}-P_{E^{\prime}}\|_{{\rm{opr}}}\leq\|P_{\mathcal{E}}-P_{{\mathcal{E}}^{\prime}}\|_{{\rm{opr}}}+2\|V-V^{\prime}\|_{{\rm{opr}}}.$
Combining (4.4), (4.3) and Lemma 4.7, we obtain that
$\|P_{E}-P_{E^{\prime}}\|_{\rm{opr}}\leq\frac{2^{5/2}d^{1/2}\epsilon n\|A_{\mu}\|_{{\rm{opr}}}}{\gamma(\lambda)}+2\sqrt{\frac{3n\epsilon}{\mu_{\min}}},$
which finishes the proof.
Theorem 4.6 shows that the difference in the Fourier transform is small if $\epsilon$, the bound on the relative perturbation of the block sizes, is much smaller than both $\frac{\mu_{\min}}{n}$ and $\frac{\gamma(\lambda)}{nd^{1/2}\|A_{\mu}\|_{{\rm{opr}}}}$. In particular, in the setting of a ${\mathbb{G}}$-SBM with almost uniform block sizes, Theorem 4.6 gives the conditions under which the well-structured group Fourier basis obtained in Theorem 4.4 can be used as a good approximation for the ${\mathbb{G}}$-SBM Fourier basis. In the next section, we study ${\mathbb{G}}$-SBM with block sizes that are not close to uniform.
4.3 ${\mathbb{G}}$-SBM with general measure
For samples of a ${\mathbb{G}}$-SBM with non-uniform measure, the character basis described in Section 4.1 is not necessarily a good choice. Namely, this basis does not provide convergence of the graph Fourier transform as described in Theorem 2.9. In this section we discuss how the group symmetries of the Cayley matrix help to partially define a graphon Fourier basis (Theorem 4.10). First, we introduce the necessary notation.
Notation 1.
We define the index identifier function $\iota$ as follows. Fix the ordering $\widehat{{\mathbb{G}}}=\{\chi_{1},\ldots,\chi_{n}\}$, and assume that $\chi_{1}$ is the trivial representation (i.e., $\chi_{1}\equiv 1$). For every $\chi\in\widehat{{\mathbb{G}}}$, let $\iota(\chi)$ denote the index $i$ such that $\chi=\chi_{i}$.
The next theorem reduces the problem of computing an eigenbasis for the model matrix $W$ to the same problem for the smaller $n\times n$ matrix $\widetilde{M}\Gamma$. Both $\widetilde{M}$ and $\Gamma$ are derived from the characters of $A$ and are easy to compute. This reduction may give us insight into properties of the eigenvectors of $W$ that might not be obvious from $A_{\mu}$ (e.g., see Proposition 4.12).
Theorem 4.10 (GFT for samples of ${\mathbb{G}}$-SBM with general measure).
Let $W$ be the model matrix of ${\rm{SBM}}(A,\mu,N)$, where $A$ is a Cayley matrix on an Abelian group ${\mathbb{G}}$. Define
$\widetilde{M}=\left[\frac{1}{n}\sum_{j=1}^{n}\mu_{j}(\chi_{k}^{-1}\chi_{\ell})(g_{j})\right]_{k,\ell\in[n]}.$
Let $U,\Gamma$ be so that $A=U\Gamma U^{*}$ is the diagonalization of $A$ given in Lemma 4.2, and let $D$ be as in Definition 3.1. For a unit vector $Y\in{\mathbb{C}}^{N}$ and $\lambda\neq 0$, we have
$Y\ \mbox{ is a }\ \lambda\mbox{-eigenvector of }W\iff\ \frac{1}{\sqrt{N}}(U^{*}D^{\rm{T}})Y\ \mbox{ is a }\frac{\lambda}{N}\mbox{-eigenvector of }\ \widetilde{M}\Gamma.$
Moreover, $\widetilde{M}$ satisfies certain symmetries, in the sense that $\widetilde{M}_{k,\ell}=\widetilde{M}_{1,\iota(\chi_{k}^{-1}\chi_{\ell})}$ for $\chi_{k},\chi_{\ell}\in\widehat{{\mathbb{G}}}$.
Proof 4.11.
Let $Y\in{\mathbb{C}}^{N}$, $Z\in{\mathbb{C}}^{n}$, and $\lambda\neq 0$. With notation as in Definition 3.1, we have
(4.5) $Y=\frac{1}{\sqrt{N}}(DM^{-1}U)Z\iff Z=\frac{1}{\sqrt{N}}(U^{*}D^{\rm{T}})Y\ \mbox{ and }\ Y\in{\rm range}(D).$
In fact, Eq. 4.5 can be easily verified using the equation $\frac{1}{N}D^{\rm{T}}D=M$ (Lemma 3.2 Item (a)). Moreover, if $Y,Z$ satisfy the relation described in (4.5), then $Y$ is nonzero precisely when $Z$ is nonzero. This follows from the fact that $DM^{-1}U$ and $U^{*}D^{\rm{T}}D=N(U^{*}M)$ have null kernels.
Next note that putting Item (a) and Item (b) of Proposition 3.4 together, for $\lambda\neq 0$, we have the following equivalence:
$WY=\lambda Y\quad\mbox{ \emph{iff} }\quad\exists X\in{\mathbb{C}}^{n},\ Y=VX\text{ and }A_{\mu}X=\frac{\lambda}{N}X.$
Substituting $A=U\Gamma U^{*}$ in the expression $A_{\mu}=M^{\frac{1}{2}}AM^{\frac{1}{2}}$, the equality $A_{\mu}X=\frac{\lambda}{N}X$ can be equivalently written as $\Gamma U^{*}M^{\frac{1}{2}}X=\frac{\lambda}{N}U^{*}M^{-\frac{1}{2}}X.$ Letting $Z=U^{*}M^{\frac{1}{2}}X$ (equivalently $X=M^{-\frac{1}{2}}UZ$), the above iff statement can be rewritten as
$WY=\lambda Y\quad\mbox{ \emph{iff} }\quad\exists Z\in{\mathbb{C}}^{n},\ Y=VM^{-\frac{1}{2}}UZ\text{ and }\Gamma Z=\frac{\lambda}{N}(U^{*}M^{-1}U)Z.$
From $\Gamma Z=\frac{\lambda}{N}(U^{*}M^{-1}U)Z$, we observe that $Z$ is a $\lambda$-eigenvector of $N(U^{*}MU)\Gamma$. In addition, $Y=VX=\frac{1}{\sqrt{N}}(DM^{-1}U)Z$, so $Z$ and $Y$ satisfy the equations from Eq. 4.5. So to finish the proof, we only need to show that $\widetilde{M}=U^{*}MU$.
For $k,l\in[n]$, a direct calculation shows that the $(k,l)$-entry of $U^{*}MU$ is given by
$\frac{1}{n}\sum_{j=1}^{n}\mu_{j}\overline{\chi_{k}(g_{j})}\chi_{l}(g_{j})=\frac{1}{n}\sum_{j=1}^{n}\mu_{j}(\chi_{k}^{-1}\chi_{l})(g_{j}).$
Finally, let $\chi,\chi_{k},\chi_{\ell}\in\widehat{{\mathbb{G}}}$ and suppose $\iota(\chi\chi_{k})=k^{\prime},\iota(\chi\chi_{\ell})=\ell^{\prime}$. Then, the $(k^{\prime},\ell^{\prime})$-entry of $U^{*}MU$ is given by $\frac{1}{n}\sum_{j=1}^{n}\mu_{j}\overline{\chi_{k^{\prime}}(g_{j})}\chi_{\ell^{\prime}}(g_{j})=\frac{1}{n}\sum_{j=1}^{n}\mu_{j}\overline{\chi(g_{j})\chi_{k}(g_{j})}\chi(g_{j})\chi_{\ell}(g_{j}),$ which simplifies to the $(k,\ell)$-entry of $U^{*}MU$, since $\overline{\chi(g_{j})}\chi(g_{j})=1$ for all $j$. Taking $\chi=\chi_{k}^{-1}$, and noting that $\chi_{1}$ is the identity element of the group $\widehat{{\mathbb{G}}}$ finishes the proof.
The applicability of Theorem 4.10 in practice is due to the fact that $\widetilde{M}$ has certain symmetric features, which result in simplified computations for the eigen-decomposition of $\widetilde{M}\Gamma$. We discuss the case of ${\mathbb{G}}$-SBM models with a dominant block below, where we can offer a recipe for computing eigenvectors of $\widetilde{M}\Gamma$ with zero mean.
Proposition 4.12.
Let ${\mathbb{G}}$ be an Abelian group of size $n$ with neutral element $e_{{\mathbb{G}}}$. Let $W$ be the model matrix of ${\rm{SBM}}(A,\mu,N)$, where $A$ is a Cayley matrix on ${\mathbb{G}}$, and let $V$ be the isometry as in Definition 3.1. Let $\tau\in(0,\frac{1}{n})$ be such that $\tau N$ is an integer. Define the measure $\mu$ as follows:
$\mu(\{e_{{\mathbb{G}}}\})=1-\tau(n-1)\ \mbox{ and }\ \mu(\{g\})=\tau,\forall\ g\in{\mathbb{G}}\setminus\{e_{{\mathbb{G}}}\}.$
  • (i)
    Let $m>1$. If $\gamma$ is an eigenvalue of $A$ of multiplicity $m$ then $N\tau\gamma$ is an eigenvalue of $W$ of multiplicity at least $m-1$.
  • (ii)
    Let $\{\chi_{\alpha_{i}}:1\leq i\leq m\}$ be an eigenbasis of characters for the $\gamma$-eigenspace of $A$ as in Lemma 4.2. Then $\left\{V\left((i-1)\chi_{\alpha_{i}}-\sum_{j=1}^{i-1}\chi_{\alpha_{j}}\right):1<i\leq m\right\}$ are orthogonal $N\tau\gamma$-eigenvectors of $W$.
Proof 4.13.
Let ${\mathbb{G}}=\{g_{1},\ldots,g_{n}\}$ be labeled such that $g_{1}=e_{{\mathbb{G}}}$. Let $\chi_{1}$ denote the trivial character of ${\mathbb{G}}$, i.e., $\chi_{1}(g)=1$ for all $g\in{\mathbb{G}}$. Using the orthogonality of characters, we get
$\sum_{g\in{\mathbb{G}}\setminus\{e_{{\mathbb{G}}}\}}\chi_{k}^{-1}(g)\chi_{l}(g)$ $=$ $\sum_{g\in{\mathbb{G}}\setminus\{e_{{\mathbb{G}}}\}}\overline{\chi_{k}(g)}\chi_{l}(g)=\langle\chi_{l},\chi_{k}\rangle-\overline{\chi_{k}(e_{{\mathbb{G}}})}\chi_{l}(e_{{\mathbb{G}}})=\left\{\begin{array}[]{cc}-1&k\neq l\\ n-1&k=l\end{array}\right..$
So the matrix $\widetilde{M}$ from Theorem 4.10 is given by $\widetilde{M}=\frac{1-n\tau}{n}J+\tau I$, where $J$ is the all 1 matrix, and $I$ is the identity matrix.
[…]
Using the notations from Theorem 4.10, let $Y_{i}=\frac{1}{\sqrt{N}}DM^{-1}U{\nu_{i}}$ for $1<i\leq m$. We have
$Y_{i}$ $=$ $\frac{1}{\sqrt{N}}DM^{-1}{\left((i-1)\chi_{\alpha_{i}}-\sum_{j=1}^{i-1}\chi_{\alpha_{j}}\right)}=VM^{-\frac{1}{2}}{\left((i-1)\chi_{\alpha_{i}}-\sum_{j=1}^{i-1}\chi_{\alpha_{j}}\right)}$
$=$ $V\left(\frac{1}{\sqrt{\tau}}\left(I+{\rm diag}(\sqrt{\frac{\tau}{1-(n-1)\tau}}-1,0,\ldots,0)\right)\right){\left((i-1)\chi_{\alpha_{i}}-\sum_{j=1}^{i-1}\chi_{\alpha_{j}}\right)}.$
[…]
Remark 4.14 (Alternative interpretation of Theorem 4.10).
The previous theorem can be understood in the context of weighted Fourier analysis on Abelian groups [5]. Namely, we can interpret the effect of the weight $\mu$ as a change of measure on $\widehat{{\mathbb{G}}}$. Consider $\mu$ as the natural probability measure on $\widehat{{\mathbb{G}}}$ given by $\mu(\{\chi_{i}\})=\mu_{i}$, and define the weighted inner product space $\ell^{2}(\widehat{{\mathbb{G}}},\mu)$ via $\langle f,g\rangle_{\ell^{2}(\widehat{{\mathbb{G}}},\mu)}=\sum_{\chi\in\widehat{{\mathbb{G}}}}f(\chi)\overline{g(\chi)}\mu(\chi)$. We have
(4.7) $(\widetilde{M}\Gamma)_{k,l}=\frac{1}{n}\,(\mathcal{F}f)(\chi_{l})\langle\chi_{l},\chi_{k}\rangle_{\ell^{2}(\widehat{{\mathbb{G}}},\mu)},\quad$
where $\mathcal{F}f$ denotes the classical Fourier transform of $f$ as a function on the Abelian group ${\mathbb{G}}$. To interpret the above equation, note that when the block sizes are equal, the measure $\mu$ is uniform. In this case, the orthogonality of group characters imply that the matrix $\widetilde{M}\Gamma$ is diagonal (as it should be), since $\langle\chi_{l},\chi_{k}\rangle_{\ell^{2}(\widehat{{\mathbb{G}}},\mu)}=\frac{1}{n}\langle\chi_{l},\chi_{k}\rangle_{{\mathbb{C}}^{n}}=0$ when $l\neq k$. The formula in (4.7) allows us to witness non-uniformity of block sizes in terms of a non-diagonal $\widetilde{M}\Gamma$.
5 Example: ${\mathbb{Z}}_{5}$-SBMs with different block sizes
To illustrate the theory presented in Section 3 and Section 4, we consider stochastic block models based on the cyclic group $\mathbb{Z}_{5}$ with various block sizes. The group $\mathbb{Z}_{5}$ is equipped with addition modulo 5, and its elements are conventionally represented as $\{0,1,2,3,4\}$. However, for consistency with the rest of the paper, we relabel these elements as $g_{i}=i-1$ for $1\leq i\leq 5$.
We fix $A$ to be the Cayley matrix on ${\mathbb{Z}}_{5}$ associated with the connection function $f$ defined as $f(1)=f(4)=0.8$ and $f(i)=0.2$ for $i=0,2,3$. More precisely, the probability matrix $A=[a_{i,j}]$ is given by $a_{i,j}=f(j-i)$ for $1\leq i,j\leq 5$. Let $\mu=\{\mu_{i}\}_{i=1}^{5}$ be the measure representing the block sizes of SBMs. All computations presented in this section were done using Mathematica version 13.2.1. Throughout this section, we use notations from Definition 3.1.
The set of characters $\widehat{\mathbb{Z}_{5}}=\{\chi_{1},\dots,\chi_{5}\}$ is generated under pointwise multiplication by the character $\chi:\mathbb{Z}_{5}\to\mathbb{T}$, defined as $\chi(j)=\omega^{j}$, where $\omega=e^{\frac{2\pi i}{5}}$ is a primitive 5th root of unity. More explicitly, we set $\chi_{j}=\chi^{j-1}$ for $1\leq j\leq 5$. Applying Lemma 4.2, the eigenvalues of $A$ are given by
$\lambda_{i}:=\sum_{j=0}^{4}f(j)\omega^{-(i-1)j},\quad\text{for }1\leq i\leq 5.$
Since $f(j)=f(-j)$, it follows that $\lambda_{2}=\lambda_{5}$ and $\lambda_{3}=\lambda_{4}$. As stated in Theorem 4.10, we utilize the eigen-decomposition of $\widetilde{M}\Gamma$ to determine the SBM Fourier basis for this graphon, where $\Gamma={\rm diag}(\lambda_{1},\ldots,\lambda_{5})$ and
$\widetilde{M}=\left[\frac{1}{5}\sum_{j=1}^{5}\mu_{j}(\chi_{k}^{-1}\chi_{l})(g_{j})\right]_{k,l}=\left[\frac{1}{5}\sum_{j=1}^{5}\mu_{j}\omega^{(l-k)(j-1)}\right]_{k,l}.$
5.1 Comparing the graph Fourier basis with the SBM Fourier basis
We first sampled from a ${\mathbb{Z}}_{5}$-SBM and compared the graph Fourier basis with the SBM Fourier basis. Recall that the graph Fourier basis is an eigenbasis of the graph adjacency matrix, while the SBM Fourier basis is an eigenbasis of $W$, the model matrix. We sampled from ${\rm{SBM}}(A,\mu,N)$ where $A$ is the Cayley matrix described above, $N=1200$, and $\mu_{1}=1/3$, $\mu_{2}=\dots=\mu_{5}=1/6$ (so there is one block of size $400$ and four of size $200$). We computed the ${\mathbb{Z}}_{5}$-SBM Fourier basis from the matrix $\widetilde{M}\Gamma$ as described in Theorem 4.10.
We took six different samples from the ${\rm{SBM}}(A,\mu,N)$. The samples yielded very similar results, so we include the results of one of them. Table 1 gives the eigenvalues of the model matrix as given by Theorem 4.10 compared with the first six eigenvalues of the adjacency matrix of the sampled graph. As explained in Section 3.1, since ${\rm{SBM}}(A,\mu,N)$ consists of five blocks, the SBM-driven Fourier basis only takes the eigenvectors corresponding to the first five eigenvalues into account, as the projection on the remaining eigenvectors may be viewed as noise. We include results for the sixth eigenvalue to show that indeed, $\lambda_{6}$ of the sample is significantly smaller (in magnitude) than the other eigenvalues. It is apparent from the table that the nonzero eigenvalues of $W$ have multiplicity 1, so we have a unique eigenbasis for the range of $W$.
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $\lambda_{5}$ $\lambda_{6}$
Model matrix 2622.1 -1290.3 -970.82 468.1 370.8 0
Sample adjacency matrix $2$1291.6 -970.85 469.8 373.1 -61
Table 1: Eigenvalues of the model matrix $W$ compared with eigenvalues of the adjacency matrix of a sample.
We then compared the first five eigenvectors $\phi_{i}$ of the model matrix $W$ of the SBM with the corresponding eigenvectors $\varphi_{i}$ of the adjacency matrix of the sampled graph. Our results show that the inner product between each pair of eigenvectors is at most $10^{-4}$, implying that the angle between the corresponding eigenvectors is no greater than $10^{-2}$. This close agreement aligns with the convergence result stated in Theorem 2.9.
5.2 Comparing a transferred character basis with the SBM-basis
Let $A$ be the Cayley matrix as given in Section 5.1. Using the characters of ${\mathbb{Z}}_{5}$, we fix the eigenbasis of $A$ consisting of the following real-valued unit vectors:
$\phi_{1}=\frac{\chi_{1}}{\sqrt{5}},\,\phi_{2}=\frac{\chi_{3}+\chi_{4}}{\sqrt{10}},\,\phi_{3}=\frac{\chi_{4}-\chi_{3}}{i\sqrt{10}},\,\phi_{4}=\frac{\chi_{2}+\chi_{5}}{\sqrt{10}},\,\phi_{5}=\frac{\chi_{2}-\chi_{5}}{i\sqrt{10}}.$
For $1\leq k\leq 20$, define the measure $\mu^{k}$ on ${\mathbb{Z}}_{5}$ to be
$\mu^{k}_{1}=\frac{60+4k}{300}\ \mbox{ and }\ \mu^{k}_{i}=\frac{60-k}{300}\ \mbox{ for }2\leq i\leq 5.$
We set $N=3000$, and consider the isometry $V_{k}$ associated with ${\rm{SBM}}(A,\mu^{k},N)$ given in Definition 3.1 (d). Inspired by Proposition 3.4 (b), we use $V_{k}$ to transfer the basis $\{\phi_{i}\}_{i=1}^{5}$ of $A$ to an orthonormal set $\{\xi_{i}^{k}\}_{i=1}^{5}\subseteq{\mathbb{C}}^{N}$ defined as $\xi_{i}^{k}:=V_{k}\phi_{i}$. We call $\{\xi_{i}^{k}\}_{i=1}^{5}$ the transferred character basis for the range of the model matrix $W_{k}$ of ${\rm{SBM}}(A,\mu^{k},N)$. Note that the transferred character basis is not an eigenbasis of the range of $W_{k}$ unless $\mu^{k}$ is uniform. We now present a numerical study to compare the transferred character basis with the SBM-basis $\{Y_{i}^{k}\}_{i=1}^{n}$ for $W_{k}$ derived in Theorem 4.10.
This provides us with a class of examples of SBMs in which one block is larger than the rest, and the difference in block sizes becomes more pronounced as $k$ increases.
(a) Plot of the inner products between transferred character basis $\{\xi_{i}^{k}\}_{i=1}^{5}$ and the SBM Fourier basis $\{Y^{k}_{i}\}_{i=1}^{5}$ for $k=1,\ldots,20$
(b) Limiting graphon when $k=0$
(c) Limiting graphon when $k=20$
Figure 2: Comparison of transferred character basis with SBM Fourier basis for non-uniform measures
Fig. 2 gives the inner product of the corresponding eigenvectors in the two bases. That is, for a given $i$ and $k$ the inner product $\langle\xi_{i}^{k},Y_{i}^{k}\rangle$ is displayed. We see that the agreement between the transferred character basis and the SBM Fourier basis is quite good, although it does decrease as the block sizes become less uniform. Fig. 2 also demonstrates Proposition 4.12, since we are in the case of one large block. By this proposition, $\xi_{3}^{k}$ and $\xi_{5}^{k}$ are eigenvectors of $W_{k}$, and thus belong to the SBM Fourier basis. We see in the figure that indeed $\xi_{3}^{k}$ and $Y_{3}^{k}$ are in perfect agreement for all $k$. The same is true for $\xi_{5}^{k}$ and $Y_{5}^{k}$.
5.3 Three different block sizes
Next, we consider the measure $\mu^{k}_{1}=\frac{30+2k}{150}$, $\mu^{k}_{2}=\frac{30+k}{150}$ and $\mu^{k}_{3}=\mu^{k}_{4}=\mu^{k}_{5}=\frac{30-k}{150}$ and take $N=150$. We let $k$ vary from 0 to 20. Again, we compared the transferred character basis, computed as in the previous section, with the SBM Fourier basis, computed using Theorem 4.10. For each $k$, let $\{Y^{k}_{i}\}_{i=1}^{N}$ be the eigenvectors of the model matrix $W_{k}$.
As before, we compute the inner product between $\{Y^{k}_{i}\}_{i=1}^{5}$ and the corresponding vectors of the transferred character basis $\{\xi^{k}_{i}\}_{i=1}^{5}$. The results are shown in Fig. 3 (a). At $k=0$, all block sizes are equal so we have perfect agreement. As $k$ increases, the agreement deteriorates more quickly than in the case of just two distinct block sizes.
(a) Plot of the magnitude of the inner product between the transferred character basis $\{\xi^{k}_{i}\}_{i=1}^{5}$ and the SBM Fourier basis $\{Y^{k}_{i}\}_{i=1}^{5}$ for $k=0,\ldots,20$.
(b) Inner product between the transferred character basis elements $\{\xi^{m}_{i}\}_{i=1}^{5}$ and the corresponding SBM basis $\{Y^{m}_{i}\}_{i=1}^{5}$ of model 1 from Section 5.2 (blue) and model 2 from Section 5.3 (red).
Figure 3: Comparison of graphon with 1 perturbed block versus 2 perturbed blocks
In Fig. 3 (b), we compare two different models. Model 1 is constructed from a measure that gives blocks of size 2000, 250, 250, 250, 250, and Model 2 is constructed from blocks of size 1350, 1344, 1102, 1102, 1102. We observe that the agreement between the transferred character basis and the actual SBM basis for model 2 is strong. However, for model 1 there is only strong agreement for $\xi_{3}$ and $\xi_{5}$. This is despite the relatively large perturbations in the block sizes present in model 1. As we saw earlier, $\xi_{3}$ and $\xi_{5}$ are SBM basis elements in model 1. Model 2 has smaller deviation from the uniform measure, and gives better performance overall.
Acknowledgements
MG acknowledges support from NSF Grant DMS-2408008 and Simons Travel Support for Mathematicians. JJ was funded by an NSERC Discovery Grant. This paper was initiated while two of the authors conducted an RIT visit at the Banff International Research Center. We are grateful to BIRS for the financial support and hospitality.
Appendix A Implementing the SBM Fourier transform
Let $G$ be a graph sampled from ${\rm{SBM}}(A,\mu,N)$ and let $X$ be a signal on $G$. In this section, we show exactly how to compute the SBM-driven Fourier transform of $X$; a summary of the steps is given in Fig. 4.
Let $A_{\mu}=U_{0}\Lambda U_{0}^{*}$ be a spectral decomposition of $A_{\mu}$, where $U_{0}$ is a unitary matrix and $\Lambda$ is an $n\times n$ diagonal matrix containing the eigenvalues of $A_{\mu}$. We first remove from $U_{0}$ all columns corresponding to eigenvalue zero; we denote the new matrix by $U_{0}$ again. The remaining columns of $U_{0}$ span $\ker(A_{\mu})^{\perp}$, the orthogonal complement to the kernel of $A_{\mu}$.
1. Compute $A_{\mu}=[\sqrt{\mu_{i}\mu_{j}}a_{i,j}]_{i,j}$. 2. Find the spectral decomposition $A_{\mu}=U_{0}\Lambda U_{0}^{*}$. 3. Remove all columns from $U_{0}$ that correspond to zero eigenvalues of $A_{\mu}$. 4. Set $k_{i}=\mu_{i}N$. Compute $V=[v_{i,j}]_{N\times n}$ as $v_{i,j}=\left\{\begin{array}[]{ll}\frac{1}{\sqrt{k_{j}}}&\mbox{ if }\ \sum_{s=1}^{j-1}k_{s}<i\leq\sum_{s=1}^{j}k_{s}\\ 0&\mbox{ otherwise}\end{array}\right.$ 5. Compute $U=VU_{0}$. 6. The SBM-driven Fourier transform is defined as follows: (a) For each distinct eigenvalue $\lambda$, the Fourier projection is $\widehat{X}(\lambda)=\sum_{i:\ \lambda_{i}=\lambda}\langle X,U_{i}\rangle U_{i}.$ (b) Compute $\widehat{X}(0)=X-\sum_{\lambda\not=0}\widehat{X}(\lambda).$ (c) The inverse Fourier transform is $X=\sum_{\lambda}\widehat{X}(\lambda)$.
Figure 4: Algorithm to compute the SBM-driven Fourier transform of a signal $X$ on a graph sampled from ${\rm{SBM}}(A,\mu,N)$.
If all nonzero eigenvalues of $A_{\mu}$ are distinct then we can adopt a simplified approach. In this case, all eigenspaces corresponding to nonzero eigenvalues are 1-dimensional, so we can represent the Fourier coefficients as scalars. Let $U=VU_{0}$, where $V$ is the isometry given in Definition 3.1 (d). By Proposition 3.4, the columns of $U$ are exactly the eigenvectors of $W$ corresponding to its nonzero eigenvalues, and they form an orthonormal basis for $\ker(W)^{\perp}$. The SBM-driven Fourier transform of the graph signal $X\in{\mathbb{C}}^{N}$ can then be computed as $\widehat{X}=U^{*}X$. In particular, the Fourier coefficient associated with non-zero eigenvalue $\lambda_{i}$ equals the inner product $\widehat{X}(\lambda_{i})=\langle X,U_{i}\rangle$, where $U_{i}$ is the eigenvector of $W$ associated with eigenvalue $N\lambda_{i}$. The Fourier coefficient $\widehat{X}_{0}$ associated with eigenvalue 0 can be computed as $\widehat{X}_{0}=\|P_{0}(X)\|=\|X-\sum_{i}\langle X,U_{i}\rangle U_{i}\|$. The (partial) inverse transform can be obtained as $\widetilde{X}=U\widehat{X}$. If $X\in\ker(W)^{\perp}$, then $\widehat{X}_{0}=0$ and $\widetilde{X}=X$. Otherwise, $\widetilde{X}=X-P_{0}(X)$.
In the general case where $A_{\mu}$ has repeated eigenvalues, we cannot use the individual inner products $\langle X,U_{i}\rangle$ to represent the Fourier coefficients since they do not have the required convergence properties (see Section 2.2). In this case, we need to adopt the graphon Fourier transform described in Definition 2.6. Adapted to graph signals in ${\mathbb{C}}^{N}$, the Fourier projection corresponding to nonzero eigenvalue $\lambda$ of $A_{\mu}$ is represented by the vector
$\widehat{X}(\lambda)=\sum_{i:\lambda_{i}=\lambda}\langle X,U_{i}\rangle U_{i}.$
For eigenvalue zero, we have that $\widehat{X}(0)=X-\sum_{i:\lambda_{i}\neq 0}\widehat{X}(\lambda)$. The inverse Fourier transform of $X$ is given as $X=\sum_{\lambda}\widehat{X}(\lambda)$, where the sum is over all eigenvalues of $A_{\mu}$, including zero.
Appendix B Matrix perturbation theory
The relation between the eigenvalues and eigenvectors of a symmetric, real-valued matrix $A$ and a perturbed matrix $\tilde{A}$ has long been the object of study. For an overview, see [38]. In the following, let $A$ and $\tilde{A}$ be symmetric matrices in ${\mathbb{R}}^{n\times n}$, where $H=A-\tilde{A}$ is considered to be the error between the matrix $A$ and its perturbation $\tilde{A}$. Assume that $A$ has eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{n}$, and $\tilde{A}$ has eigenvalues $\tilde{\lambda}_{1}\geq\tilde{\lambda}_{2}\geq\dots\geq\tilde{\lambda}_{n}$. Where appropriate, we use ‘dummy values’ $\lambda_{0}=\infty$ and $\lambda_{n+1}=-\infty$.
The difference between eigenspaces of $A$ and $\tilde{A}$ can be bounded via the Davis-Kahan theorem, which is stated in terms of principal angles between subspaces. Let ${\mathcal{E}}$ and $\tilde{{\mathcal{E}}}$ be $d$-dimensional subspaces of ${\mathbb{R}}^{n}$, and let $V$ and $\tilde{V}$ be $n\times d$ matrices whose columns form an orthonormal basis of ${\mathcal{E}}$ and $\tilde{{\mathcal{E}}}$, respectively. The principal angles of $({\mathcal{E}},\tilde{{\mathcal{E}}})$ are the angles $\theta_{i}=\cos^{-1}(\rho_{i})$, $i\in[d]$, where $\rho_{1},\rho_{2},\dots,\rho_{d}$ are singular values of $V^{\rm{T}}\tilde{V}$. Let $\Theta={\rm diag}(\theta_{1},\theta_{2},\dots,\theta_{d})$ be the matrix representing the principal angles, and let $\sin(\Theta)$ be the matrix obtained from $\Theta$ by taking the sine component-wise. The distance between ${\mathcal{E}}$ and $\tilde{{\mathcal{E}}}$ is given by
$\|\sin(\Theta({\mathcal{E}},\tilde{{\mathcal{E}}}))\|_{F},$
where $\|\cdot\|_{F}$ denotes the Frobenius norm. We now give a bound for this distance. The bound is taken from [39] and is a variant of the Davis-Kahan theorem. This bound only requires the gap in eigenvalues of one of the matrices, making it suitable for a context where one of the matrices has known properties, and the second matrix is a perturbation of the first.
Theorem B.1 ([39], Davis-Kahan).
Let $A$ and $\tilde{A}$ be two symmetric matrices in ${\mathbb{R}}^{n\times n}$, with eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{n}$ and $\lambda^{\prime}_{1}\geq\lambda_{2}^{\prime}\geq\dots\geq\lambda^{\prime}_{n}$, respectively. Fix $1\leq r\leq s\leq n$ and assume that $\lambda_{r}=\lambda_{r+1}=\dots=\lambda_{s}=\lambda$. Assume $\min\{\lambda_{r-1}-\lambda,\lambda-\lambda_{s+1}\}>0$. Let $d=s-r+1$ be the multiplicity of $\lambda$. Let ${\mathcal{E}}$ be the eigenspace of $A$ corresponding to $\lambda$, and let $\tilde{{\mathcal{E}}}$ be the space spanned by orthogonal eigenvectors of $\tilde{A}$ corresponding to eigenvalues $\tilde{\lambda}_{r},\tilde{\lambda}_{r+1},\dots,\tilde{\lambda}_{s}$. Then
$\|\sin(\Theta({\mathcal{E}},\tilde{{\mathcal{E}}}))\|_{F}\leq\frac{2\min\{d^{1/2}\|A-\tilde{A}\|_{{\rm{opr}}},\|A-\tilde{A}\|_{F}\}}{\min\{\lambda_{r-1}-\lambda,\lambda-\lambda_{s+1}\}}$
An alternative way of representing the distance between the subspaces, which is relevant in the context of the graph and graphon Fourier transforms, is by quantifying the difference in the projection operators on the two subspaces. Precisely, let $P_{{\mathcal{E}}}$ and $P_{\tilde{{\mathcal{E}}}}$ be the projection operators onto ${\mathcal{E}}$ and $\tilde{{\mathcal{E}}}$, respectively. Note that $P_{{\mathcal{E}}}$ and $P_{\tilde{{\mathcal{E}}}}$ are represented by the matrices $VV^{\rm{T}}$ and $\tilde{V}{\tilde{V}}^{\rm{T}}$, respectively. We can represent the distance between ${\mathcal{E}}$ and $\tilde{{\mathcal{E}}}$ in terms of the Frobenius distance between these matrices.
Let $\Theta=\Theta({\mathcal{E}},\tilde{{\mathcal{E}}})$. Then $\cos(\Theta)$ is the diagonal matrix containing the singular values of $V^{\rm{T}}\tilde{V}$. Using the singular value decomposition, we have that there exist orthogonal matrices $U_{1},U_{2}$ so that $V^{\rm{T}}\tilde{V}=U_{1}\cos(\Theta)U_{2}^{\rm{T}}$. Now let $W\in{\mathbb{R}}^{n\times n-d}$ be the matrix whose columns form an orthonormal basis of the space orthogonal to $\tilde{{\mathcal{E}}}$. Thus $WW^{\rm{T}}=I-\tilde{V}\tilde{V}^{\rm{T}}$ is the projection onto $\tilde{{\mathcal{E}}}^{\perp}$. We then have that
$V^{\rm{T}}WW^{\rm{T}}V+V^{\rm{T}}\tilde{V}\tilde{V}^{\rm{T}}V=V^{\rm{T}}V=I_{d},$
since $V$ is orthogonal. Now $V^{\rm{T}}\tilde{V}(\tilde{V}^{\rm{T}}V)=U_{1}\cos(\Theta)^{2}U_{1}^{\rm{T}}$ and thus
$(V^{\rm{T}}W)(V^{\rm{T}}W)^{\rm{T}}=I_{d}-U_{1}\cos(\Theta)^{2}U_{1}^{\rm{T}}=U_{1}\sin^{2}(\Theta)U_{1}^{\rm{T}}.$
So we get
$\|\sin(\Theta({\mathcal{E}},\tilde{{\mathcal{E}}}))\|^{2}_{F}=\|V^{T}W\|_{F}^{2}=\|P_{{\mathcal{E}}}(I-P_{\tilde{{\mathcal{E}}}})\|^{2}_{F}.$
Similarly, we have that $\|\sin(\Theta({\mathcal{E}},\tilde{{\mathcal{E}}}))\|^{2}_{F}=\|(I-P_{{\mathcal{E}}})P_{\tilde{{\mathcal{E}}}}\|^{2}_{F}$. Putting these two expressions together with the Davis-Kahan theorem we obtain the following corollary.
Corollary B.2.
(B.1) $\|P_{\mathcal{E}}-P_{\tilde{{\mathcal{E}}}}\|_{F}\leq\frac{2^{3/2}\min\{d^{1/2}\|A-\tilde{A}\|_{{\rm{opr}}},\|A-\tilde{A}\|_{F}\}}{\min\{\lambda_{r-1}-\lambda,\lambda-\lambda_{s+1}\}}.$